State the time-dependent Schrodinger equation in three dimensions. State the meaning of all the terms in the equation. 


1
Expert's answer
2020-07-21T12:36:50-0400

The time-dependent Schrödinger equation reads


"i \hbar \frac{\partial}{\partial t}\Psi\\ =-\frac{ \hbar^2}{2m}\left(\dfrac{\partial^2 \Psi}{\partial x^2}+\dfrac{\partial^2 \Psi}{\partial y^2}+\dfrac{\partial^2 \Psi}{\partial z^2}\right)+V(x,y,z)\Psi"

The function Ψ varies with time t as well as with position x, y, z.

V is assumed to be a real function and represents the potential energy of the system.